### v2

# Towards Embodied Agentic AI: Review and Classification of LLM- and VLM-Driven Robot Autonomy and Interaction

Abstract
Foundation models, including large language models (LLMs) and vision-language models (VLMs), have recently enabled novel approaches to robot autonomy and human-robot interfaces. In parallel, vision-language-action models (VLAs) or large behavior models (LBMs) are increasing the dexterity and capabilities of robotic systems. This survey paper focuses on those works advancing towards agentic applications and architectures. This includes initial efforts exploring GPT-style interfaces to tooling, as well as more complex system where AI agents are coordinators, planners, perception actors, or generalist interfaces. Such agentic architectures allow robots to reason over natural language instructions, invoke APIs, plan task sequences, or assist in operations and diagnostics. In addition to peer-reviewed research, due to the fast-evolving nature of the field, we highlight and include community-driven projects, ROS packages, and industrial frameworks that show emerging trends. We propose a taxonomy for classifying model integration approaches and present a comparative analysis of the role that agents play in different solutions in today’s literature.
{strip}
Figure 1: Illustration of different approaches to the integration of LLMs, VLMs and other foundation models into robotic systems. This survey focuses on the integration perspective and the role of such models within the wider robotic system where it is deployed. Our goal is to discuss the current state-of-the-art within the context of Embodied Agentic AI applications.
I Introduction
Large language models (LLMs) and vision-language models (VLMs) have recently enabled new modalities of interaction and reasoning in robotics. While early work focused on end-to-end pipelines, mapping raw sensory input and natural language directly to robot actions, there is growing interest in modular systems that use foundation models as high-level agents. These agents interpret user intent, generate plans, invoke robot APIs, or interface with middleware such as the Robot Operating System (ROS), without replacing the underlying robot stack.
This agentic approach offers several advantages. By integrating with existing functionality, LLMs can extend the flexibility and usability of robots without discarding tested software modules. Agents can reason over unstructured instructions, select or sequence robot capabilities, and adapt responses through interaction. Recent examples include ROS interfaces for LLMs, protocol-based control layers, and agents that coordinate multiple subsystems or external tools.
Several surveys have addressed the broader intersection of foundation models and robotics [1, 2, 3, 4, 5], with a focus on multimodal architectures and their components and primarily emphasizing the general use of multimodal models in robotic decision-making and high-level planning, as well as end-to-end learning frameworks for low-level control in domain-specific settings such as manipulation. However, with the very recent rise of emergence of embodied and generalist AI agents, no prior survey to date has examined the emerging design patterns in which AI agents interface with existing control software, libraries, or middleware. Moreover, many practical systems (GitHub-hosted projects, ROS packages, or startup prototypes) remain underrepresented in the literature despite their increasing real-world relevance and impact.
This survey addresses that gap. We review recent academic and community-driven work on AI agents in robotics, emphasizing architectures that position LLMs and VLMs as intelligent intermediaries rather than direct policy generators. Our contributions are twofold:
  • We discuss the concept of Agentic AI in robotics and distinguish it from end-to-end learning or classical symbolic planning approaches.
  • We propose a taxonomy across two dimensions: integration approach to foundation models (Section III, see also Fig. 1), and agent role (Section IV, see also Fig. 3).
Together, these elements provide an alternative view, more technical and application-oriented, to the use of foundation models in robot systems.
[…]
Figure 2: Progression of projects at the intersection of Generative AI and Robotics with a focus on LLMs, VLMs and VLAs.
II Towards Embodied Agentic AI
The concept of an embodied robotic agent was introduced in the late 1980s, grounded in the idea that intelligent behavior emerges from a system’s physical embodiment and its continuous sensorimotor interaction with the environment [17]. This laid the foundation for early end-to-end neural models that directly mapped perception to action, enabling autonomous control without explicit symbolic reasoning [18, 19, 20]. The emergence of LLMs has significantly expanded the functional scope of autonomous agents, shifting them from narrowly scoped, rule-driven systems to more flexible entities capable of generalization, contextual reasoning, and adaptive behavior [21]. At pre-ChatGPT era, Code as Policies (CaP) [6] introduced robot-executable language model generated programs (LMPs), enabling reactive and vision-based control via indirect tool calling through code generation.
The release of OpenAI’s ChatGPT in November 2022 marked a major progression in the trajectory of embodied agentic AI, catalyzing the widespread adoption of LLMs as foundational components in robotic systems. While ROS has long been a standard framework for robot software integration, initial efforts to operationalize LLMs as autonomous agents predominantly targeted ROS-compatible platforms. By 2023, early prototypes began to explicitly integrate LLMs with robotic frameworks, mainly focusing on protocol integration and bidirectional communication interfaces between LLMs and ROS.
Early exploration in this domain focused on natural language interfaces for ROS 2, including ros2ai [22], an extension to augment the ROS 2 command line interface with LLM, and ROScribe [23], a tool using LLMs to generate ROS codebases from a high-level description (e.g. a user describe desired ROS graph (sensors, topics, and functionality) and ROScribe will output Python code for ROS nodes and launch files to implement that spec). Furthermore, The ROS-LLM framework [24] was proposed to enable embodied intelligence applications. It provides a template to define your robot’s API in a config file, then handles the prompt engineering and calls to the OpenAI API, effectively acting like a customizable agent. Concurrently, Microsoft’s ChatGPT for Robotics [7] introduced a modular approach to LLM-based robot control and programming, defined a high-level API specific to each robot, leveraging prompt engineering and pre-defined function libraries to enable natural language control across diverse tasks and simulation environments. During this phase, most implementations adopted a chatbot-style design, deploying LLMs to interpret free-form textual inputs and generate appropriate commands for general robot integration.
Building on these advances, 2024 marked a transformative shift with the development of specialized LLM interfaces for robotics and VLA systems, unlocked unprecedented generalization and semantic reasoning capabilities in robotic control systems. These frameworks enabled agents to process open-ended natural language instructions, ground them in visual perception, and execute corresponding actions [14]. A notable breakthrough was RT-2 [8], which introduced a novel VLA framework by co-fine-tuning pretrained VLMs on both robotic trajectory data and internet-scale vision-language tasks. $\pi$[12] advanced the field by introducing real-time, continuous control via flow-matched diffusion policies, unifying perception, reasoning, and motion generating in a single differentiable framework.
Simultaneously, agentic middleware frameworks such as ROSA (Robot Operating System Agent) [9], RAI (Robotic AI Agent) [10], and BUMBLE [11] emerged, bridging the gap between LLM-based reasoning and traditional robotic middleware, in the pursuit of embodied AI. Initially open-sourced in late 2023, ROSA represents a significant advancement in AI-driven robotic control by implementing an LLM agent built on the LangChain framework and the ReAct (Reasoning and Acting) agent paradigm. Its core innovation was the abstraction of ROS operations into tool-enabled Python functions - including listing available nodes, reading sensor data, moving a joint, etc. This allowed ROSA to translate natural language commands into validated robot actions. ROSA also embedded safety mechanisms like parameter validation, constraint enforcement, and optional human approval for critical actions. The authors demonstrate ROSA across heterogeneous platforms, including JPL’s NeBula-Spot quadruped and NVIDIA Isaac Sim environments. However, its architecture remains tightly coupled to ROS-based middleware, limiting interoperability with non-ROS systems and constraining deployment in broader contexts. RAI is another approach in this field as a flexible embodied multi-agent framework designed to integrate LLM reasoning with robotic systems (e.g. ROS 2). Its architecture introduces a novel distributed paradigm where specialized LLM agents, including perception modules, task planners, motion controllers, and safety monitors, collaborate through well-defined roles to enable concurrent, real-time task execution while maintaining system safety and reliability. The framework’s core components (Agents, Connectors, and Tools) provide integrated capabilies for multimodel sensing and actuation, retrieval‑augmented generation through vector stores and pre‑configured agent types (voice, conversational, state‑based). The framework demonstrates particular advancement in dynamic environments, featuring capabilities like online replanning and failure recovery. Proven on both physical (Husarion ROSBot XL) and simulated (tractor and manipulator) platforms. However, the framework is constrained by the limited spatial reasoning and unreliable self-correction of LLM-based agents, which repeatedly led to failures in understanding physical boundaries, object interactions, and task completion. BUMBLE (BUilding-wide MoBiLE Manipulation) is another recent contribution that uses a unified VLM-based framework for building-wide mobile manipulation. It integrates open-world perception, dual-layered memory systems, and a wide range of motor skills, allowing effective operation over extended spatial and temporal horizons. Unlike frameworks such as ROSA, which rely heavily on ROS-based middleware integration, BUMBLE features tighter coupling with navigation systems through the use of predefined landmark images. While this provides efficient and precise navigation, it also introduces limitations, such as requiring prior manual collection of landmark imagery. Nevertheless, BUMBLE’s combination of perception and action capabilities presents a significant advance in addressing the complexity of real-world environment.
Another notable development is the concept of an MCP (Model Context Protocol) server, driving agentic use and quick integration of new capabilities into existing tools (e.g., Anthropic’s Claude or AI-driven code editors like Cursor). Their penetration in robotics is so far community-driven, with several projects openly available on GitHub. The ROS-MCP library [25], for example, connects to AI assistants such as Claude and enables interaction with installed ROS 2 applications. The capabilities include (i) launching UI applications, including the Gazebo simulator and the RViz visualizer; (ii) ROS resource management (topics, nodes); (iii) ROS system interaction (publishing to topics, calling services, send action goals); (iv) environment debugging; and (v) process management (e.g., cleaning up running ROS 2 processes). This represents a paradigm shift from other frameworks that provide a complete application system such as ROSA and RAI, with MCP servers taking a plugin-based approach instead, requiring standard AI assistant applications. Another example is the ros-mcp-server [25], which uses rosbridge to connect to either ROS 1 or ROS 2 systems through WebSockets, and contains specific interfaces to odometry and velocity control topics. A benefit of leveraging assistants like Claude is that they already have other built-in tooling, such as web search, and the ability to run code internally in a sandboxed environment. This opens the possibility to more advanced behaviors without additional implementation.
A final and more recent open-source framework for agentic AI in robotics is driven by the startup OpenMind [16]. OpenMind’s OM1 introduces a modular, hardware-agnostic AI runtime intended for various robotic platforms. Compared to earlier frameworks such as ROSA and RAI, which primarily target integration with ROS-based middleware, OM1 differentiates itself with its decentralized FABRIC coordination protocol. FABRIC enables secure identity management and interoperability among heterogeneous robotic systems, diverging from the centralized orchestration common in earlier systems. While OM1 utilizes Python-based configurations and incorporates perception-action loops similar to VLA-driven systems like RT-2, it notably emphasizes decentralization. Overall, OM1 represents a step toward more collaborative and flexible robotic ecosystems.
In summary, the past three years have seen a rapid proliferation of interfaces and integration approaches for foundation models and robotics frameworks. From ROSA’s polished LangChain tools, to community-driven MCP servers, to bespoke industry solutions, the trend is clear: robots are getting a higher-level “AI agent” layer that can understand human intent and manage the robot’s own software toolkit.
Figure 2 illustrates in a timeline format some of the key milestones and most representative works of the past three years. The top part of the figure focuses on agentic-oriented approaches, while the bottom part lists other relevant models that have shaped the advances of the field. A more comprehensive list of works, both academic papers and community projects, is available in Table II.
TABLE I: Selection of most representative works spanning the time since the release of ChatGPT and the different integration approaches proposed in this survey.
Project Description Year Integration Approach Toolset/ skillset Memory Multi-modal World model Open-source Adaptation
CaP [6] Robot-Centric formulation of language model programs 2022 Protocol Indirect tool through coding Variable logging LLM N.A
ros2ai [22] ROS 2 command line interface extension with LLM 2023 Protocol ROS 2 CLI LLM N.A
ROS-LLM [24] ROS framework for embodied intelligence applications 2023 Protocol ROS actions/services or code API Logging LLM N.A
ROScribe [23] Create ROS packages using LLMs 2023 Protocol/ Interface Custom tools - LLM - N.A
ChatGPT Robotics [7] High-level functions library for robotics application 2023 Protocol/ Interface Custom tools LLM Sys. prompt & tools
ROSA [9] ROS Agent, interact with ROS-based robotics systems 2024 Interface Custom tools Chat history VLM Sys. prompt & tools
RAI [10] AI agent framework for embodied AI 2024 Interface/ Framework Custom toolsets - VLM Custom tools
Embodied Agents [26] ROS 2 based framework for robot to understand models 2024 Interface Custom tools Semantic Memory VLM Sys. prompt & tools
BUMBLE [11] Framework for open-world perception system 2024 Interface Predefined skill library VLM Custom skills
AutoRT [27] Models for Orchestration of Robotic Agents 2024 Orchestration Pretrained RT policies - VLM/ VLA Pretrained policies
ROS-MCP [28, 25] Converts natural language to ROS command 2025 Protocol Custom tools - LLM - Custom tools
Gemini Robotics [13] Generalist VLA model 2025 Embedded Tool library - VLA - -
MoManip VLA [29] VLA for general mobile manipulation 2025 Embedded Pretrained policies - VLA ✓soon Pretrained policies
ARNA [30] General-Purpose Robotic Navigation via LVLM 2025 Embedded Custom toolsets Multimodal memory VLM ✓soon Custom tools
OpenMind OM1 [16] Modular AI runtime for agents and robots 2025 Interface/ Framework Custom toolsets Chat history VLM Custom tools
UROSA [31] Distributed AI Agents for Underwater Robot Autonomy 2025 Framework Custom toolsets/ On the fly code generation Logging VLM/ LLM Sys. prompts
TABLE II: Selection of agentic AI frameworks relevant to this survey but not developed for or tested on robotic systems.
Voyager [32] LLM-powered embodied lifelong learning agent 2023 Embedded Skill library + iterative prompt LLM - -
Odyssey [33] Agents with Open-World Skills 2025 Embedded Predefined skill library LLM - Predefined skill
III Approaches to LLMs/VLMs/VLAs and AI Agents in Robotics
We propose a taxonomy to classify the work to date based on the way that large language or foundation models are integrated into robotic systems. Instead of focusing on the role, use case or type of model, our objective here is to classify the integration from a conceptual and interaction point of view.
To that end, we propose the existence of, currently, four main directions in which foundation models are integrated into robotic systems: (i) protocol integration, where the focus is on using the model as a translator between protocols (e.g., user input to predefined toolset); (ii) interface integration, where the focus is to provide an interactive approach connecting user, robotic system and the environment; (iii) orchestration-oriented integration, where the model is in charge of managing resources, tools or subsystems; and (iv) direct or embedded integration, where the model acts as a perception or control policy, either end-to-end, or as a specific subsystem. These approaches are illustrated in Fig. 1. We also classify a selection of representative works based on their integration approach in Table II.
III-A Protocol Integration
The first category, and arguably the conceptually simplest way of integrating foundation models such as LLMs into a robotic system, is to use them as translators of user inputs.
For example, ros2ai [22] does this by translating a generic user prompt into one specific ROS 2 command line interface tool call. The interaction is limited, although possible, and the process is mostly unidirectional.
More recently, with the advent of the MCP protocol as the de-facto standard for tool use in LLMs/VLMs, multiple authors have proposed technical integrations [28, 25]. The available tooling can span multiple domains: interaction with graphical interfaces (e.g., launching a simulator), the previously mentioned execution of ROS 2 command line interface commands, or the connectivity with ROS communication paradigms (publishing / subscribing to specific topics, calling services, or executing actions. Indeed, the MCP protocol maps well with, e.g., the service and action interfaces in ROS.
A notable related approach is code-as-policies [6], where the LLM outputs Python code to perform the task (using a predefined API for sensorimotor actions). The code is then executed in a loop. Essentially the LLM “programs” the robot on the fly.
III-B Interface Integration
Keeping the tool calling capabilities of a protocol integration, but enhancing the interactiveness, we classify as interface integrations those approaches where either (i) the output of tools (often in terms of actions in the real world) affects future commands or additional tool calls from the model; or (ii) the focus is on the user interaction, e.g., in HRI applications.
This category also goes in line of the research towards more agentic applications, where foundation models gain further agency with respect to user commands, perform more complex tool calls, and often run in a loop with multiple iterations and multiple tool calls per iteration.
In this paradigm, user commands can stay high level, while the foundation models can decompose the command into a number of tool calls. Additionally, as in ROSA, RAI, or BUMBLE, actions or tool calls might span a certain amount of time, requiring action feedback to be given back to the user and/or the model.
The most recent and arguably complete open-source agentic system to date is OpenMind [16], a commercial effort that has been demonstrated in humanoid and quadruped robots, among others, and is defining a new AI-centric ”operating system” for intelligent robots.
III-C Orchestration-Oriented Integration
We separate interface from orchestration integration approaches in that the latter are focused on the problem of resource management, rather than the interaction with the user. These include mainly approaches where LLMs or VLMs act as planners or coordinators, often with other AI agents involved.
This classification category spans both the integration approach and the role of the model, which is described in Section IV, where we further describe planner and orchestration agent roles.
III-D Embedded or Direct Integration
This category encompasses approaches where the LLM (or a multimodal variant) directly produces either (i) actions for the robot in an end-to-end manner, or (ii) a specific output (e.g., to be used as a perception module). These are often referred to as robotic foundation models, in their end-to-end variant. These approaches are also well covered in existing review papers [1, 2, 3, 4]. Therefore, we only include in this survey those that are most relevant from a historical perspective, or have potential to advance towards more agentic systems. Those works are included in Table II.
III-E Agentic Frameworks
It is important to note that more complex systems might span multiple categories. For example, a system might encompass an interface agent for user input, an internal coordinator agent, and potentially both protocol and direct agents as subsystems managed by the coordinator. However, most works to date focus on one specific integration approach, owing to the relative low maturity of the field.
In that context, we can also consider solutions such as RAI or OpenMind as frameworks, as they also provide tooling and infrastructure for multi-agent coordination and support for specialist agents. However, the current stage of development and deployment is closer to the interface integration class.
{forest}
[…]
Figure 3: Taxonomy of Robotic Agents grouped by role and functionality.
IV Roles and Architectures of Robotic Agents with LLMs
As we surveyed recent LLM-integrated robotic systems, it became clear that many differ not in the tasks they perform, but in how they structure decision-making, control, and modularity. To capture these architectural distinctions, we group the agents according to their functional design rather than their application domain. This perspective enables a clearer comparison of how different systems reason, sequence actions, invoke tools or skills, and interface with low-level controllers. An overview of these agent types is provided in Figure 3.
IV-1 Planner Agents
In this paradigm, an LLM plans a sequence of actions for the robot, typically by selecting or sequencing discrete skills. The LLM does not directly control actuators; instead, it outputs high-level plans that are grounded and executed by lower-level controllers. This design allows the LLM to focus on reasoning and decomposition while ensuring safety and feasibility through separate modules.
A foundational example is Google’s SayCan [34], where the LLM (PaLM) generates possible next actions (e.g., “pick up can” or “move to kitchen”), and a learned value function (affordance model) evaluates each option based on the robot’s current state. The most feasible action is then executed by the appropriate skill module. Similarly, SELP [35] uses an LLM to generate symbolic task plans, followed by structured safety and efficiency filtering before execution. ConceptAgent [36] extends this idea further by incorporating a symbolic planner and a precondition-grounding module, enabling robust task decomposition and dynamic re-planning based on environment feedback.
In contrast to orchestration agents—which continuously interact with perception, memory, and tool APIs in real-time—planning agents produce a plan upfront or iteratively refine it, with limited API interaction during execution. While orchestration agents serve as active runtime managers, planning agents focus on task structuring and leave execution monitoring to other components.
IV-2 Orchestration Agents
As robotic systems become increasingly modular and skill-rich, a new class of agents has emerged where the LLM functions not as a planner for a single robot, but as an orchestrator that manages interactions between multiple skills, components, or even agents. These systems treat the LLM as a central controller that interprets high-level task instructions and delegates responsibility to subsystems or external tools capable of execution.
In this setting, the LLM does not execute physical actions itself, nor does it merely plan a single sequence. Instead, it performs decision-level control over which skill to invoke, which robot should act, or when to switch between behaviors. This architecture enables modularity, scalability, and adaptability, especially in multi-agent or high-skill-count environments.
For example, AutoRT [27] deploys an LLM to orchestrate a fleet of real-world mobile manipulators. The system uses prompt-based reasoning to map natural language instructions to specific skill invocations (e.g., grasping, opening, navigating), while integrating filtering layers to ensure safety and feasibility. Similarly, LABOR Agent [38] enables bimanual robotic manipulation by selecting among hundreds of pretrained skills. The LLM determines which skill or skill chain should be used given a complex manipulation instruction, using grounding modules to connect symbolic task descriptions to real-world motor behaviors.
In our categorization, orchestration agents primarily refer to systems that handle high-level task planning and assignment, either by coordinating multiple internal agents within a single robot [37], or by distributing tasks across multiple embodied agents (i.e., multiple robots) [39]. While they rely on predefined capabilities, the orchestration logic remains flexible, enabling adaptive responses to novel combinations of goals, environments, or agent configurations.
IV-3 Task-Specific Agents
Task-specific agents are designed to solve narrowly defined problems, which used to rely on specialized training data or domain-specific policies. Recent work, however, explores how large language and vision-language models can enhance task-specific performance by enabling zero-shot reasoning or dynamic planning for particular objectives. For instance, LGX enables an embodied agent to find novel, uniquely described objects (e.g., ”cat-shaped mug”) using language-guided exploration and vision-language grounding without task-specific training [42]. Similarly, NavGPT uses explicit reasoning over visual and linguistic cues to follow navigation instructions, demonstrating that language models can guide task-specific behavior without prior tuning on navigation data [41]. ARNA leverages an LVLM to dynamically orchestrate perception and control modules for general-purpose navigation, but it can also be framed as a task-specific agent when executing goal-directed exploration in unseen environments [30]. Similarly, BUMBLE demonstrates a highly capable agent for manipulation and navigation by integrating an LLM with a world model and structured memory, enabling it to follow user instructions across tasks.
IV-4 Model-Centric Agents
Unlike traditional VLA models that often require separate components for perception, reasoning, and action, each trained or tuned independently, model-centric agents adopt a unified architecture where a single model is responsible for grounding multimodal inputs (e.g., images, language, proprioception) and directly producing action outputs. Examples include LEO [43], which leverages a decoder-only large language model to integrate 2D egocentric vision, 3D point clouds, and text for both instruction following and physical interaction in 3D environments; RoboCat [44], which uses a goal-conditioned decision transformer to generalize across robotic embodiments and tasks through large-scale training and self-improvement; and RoboAgent [45], which achieves high data efficiency and broad task generalization through semantic augmentation and action chunking within a unified policy model.
IV-5 Generalist Agents
This category of agents reflects a shift toward foundation-style architectures for robotics, where a centralized reasoning model (often an LLM or a multimodal transformer) interfaces flexibly with grounded, executable components. Although each system may differ in how tools are created, selected, or executed, they share a common philosophy: general-purpose reasoning layered on top of task-specialized capabilities.
Generalist agents are typically built for multi-task, multi-domain operation. Their architectures are modular by design: the high-level model interprets goals, decomposes them into subtasks, and delegates execution to a set of low-level modules, which are usually predefined skills, learned action models, or custom tools. These execution modules may be static (pretrained) or dynamic (created and expanded during deployment).
For example, Voyager [32] is an open-ended generalist agent that autonomously explores new tasks in Minecraft by generating its own tools (as Python functions), evaluating them, and storing them for future use—effectively building a lifelong, self-curated skill library. Similarly, Code as Policies uses an LLM to generate executable Python policies that map observations directly to robot actions. ODYSSEY [33] likewise operates in an open-world setting, using an LLM to reason about tasks and select from a rich library of skills grounded in the environment. In the real world, RoboGPT [46] demonstrates generalist behavior by interpreting language instructions and invoking appropriate pretrained skills for manipulation and navigation, without needing task-specific retraining.
What distinguishes these agents is not just their task breadth, but their ability to generalize across tasks through modular reasoning and flexible skill integration—an architectural pattern increasingly central to the future of scalable robotic intelligence.
IV-6 Generalist Systemic Agents
Generalist Systemic Agents focus on building reusable, modular frameworks that simplify the development and orchestration of LLM-based robotic systems. Rather than targeting specific tasks, these approaches—such as ROSA, and RAI emphasize system-level design, where perception, reasoning, and action modules are cleanly separated and easily composed. Many of these agents are built upon general-purpose LLM frameworks like LangChain, enabling flexible memory, tool integration, and conversational interfaces. Notably, this category closely overlaps with the Agentic Frameworks paradigm discussed in Section III, as these systems rely on structured interfaces through which the LLM can invoke robotic capabilities.
V Conclusion
This survey reviewed the emerging field of agentic AI in robotics, where large language and vision-language models function as intelligent intermediaries between users and robotic systems. Unlike end-to-end learning approaches, agentic systems preserve the structure of traditional robotics pipelines, using LLMs to translate goals, generate plans, and invoke APIs or tools based on available capabilities.
We organized the current landscape into two key taxonomic dimensions: model integration patterns, and agent role. Across all dimensions, we included both academic literature and practical systems from the open-source community, or industrial applications. Most of the literature today can be classified under one specific area , e.g., with a focus on either protocol, interface, orchestration, or embedded integration of foundation models. However, there are already a number of works that propose more complex system that span multiple areas. Similarly, from the perspective of agent roles, model-centric agents operate as either orchestration or generalist agents. Therefore, we see significant value in combining different approaches and subsystems to achieve higher degrees of intelligence.
Embodied or Physical Agentic AI for robotics is still in an early but rapidly evolving stage. Future work will need to address challenges related to grounding, memory, safety, deployment efficiency, and evaluation. As modular agent frameworks continue to improve, we expect increased adoption across industrial and academic robotics, making intelligent interaction with complex systems more accessible, interpretable, and adaptive.